What is the given point-slope form in slope-intercept form for y-2=3(x+1)?
Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

y - 2 = 3(x + 1)
y - 2 = 3x + 3
y = 3x + 5
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by stanbon(75887)

What is the given point-slope form
in slope-intercept form for y-2=3(x+1)?
----------------
y-2 = 3x+3
y = 3x + 5
-----------------
Cheers,
Stan H.
-----------------